Calculate the pH of the solution in which 0.2M NH4Cl and 0.1M NH3 are present. The pKb of ammonia solution is 4.75.

Grade:11

1 Answers

Arun
5 years ago
Dear Abhishek
 
To solve pH of buffer solutions you use the Henderson - Hasselbalch equation for basic buffers:  
Because you have a basic buffer first solve for pOH  
pOH = pKb + log ([salt]/[base])  
You have pKb = 4.75  
You have [salt] = 0.2M 
You have [base] = 0.1M  
Substitute into equation:  
pOH = 4.75 + log (0.2/0.1)  
pOH = 4.75 + log 2  
pOH = 4.75 + 0.30  
pOH = 5.05  

pH = 14.00-pOH  
pH = 14.00-5.05  
pH = 8.95
 
Regards
Arun (askIITians Forum expert)